# Mechanistic Indicators of Understanding in Large Language Models

### v4

@@ 1  Introduction
The striking competence of Large Language Models (LLMs) raises an intriguing question: are they just mimicking human intelligence by relying on superficial statistics, or do they form internal structures that are sufficiently sophisticated and specific to sustain comparisons with human understanding? [^1] The issue of understanding in LLMs is distinct from that of understanding with LLMs (Sullivan, 2022; Beisbart, 2025), though the two are linked: as Sullivan effectively argues, whether we can understand with LLMs depends notably on how much understanding there is in LLMs.
@@
The most parsimonious explanation is that LLMs possess no understanding whatsoever. This soberly deflationary view receives succor from the surprisingly rudimentary principle upon which LLMs are built. At its foundation, an LLM is trained on a simple task: next-token prediction. The model is given a sequence of text and tasked with predicting the most probable next word or sub-word unit (“token”). According to the deflationary view, LLMs become good at this task by relying entirely on superficial statistical patterns within the distribution of words in the training data. This suggests that LLMs still operate fundamentally like traditional $n$-gram models, only at an unprecedented scale: by tallying how frequently different sequences of $n$ words co-occur (e.g. how often “The best pet is a” is followed by “dog” versus “cat,” etc.), they compute the probability of a word following a given sequence of “n1n-1” words. On this view, LLMs are doing nothing more than learning and then matching statistical patterns derived from extensive word-counting; they lack any grasp of the concepts, facts, and principles underlying these patterns. [^2] See, e.g., Bender and Koller (2020); Bender et al. (2021); Mondorf and Plank (2024); Merrill et al. (2021).
@@
This picture effectively casts LLMs as real-world approximations of a “Blockhead”—Ned Block’s (1981) hypothetical system that mimics intelligence by memorizing a massive table of input–output pairs. [^3] Millière and Buckner (2024) discuss the relationship between LLMs and Blockhead in detail. While LLMs may be more flexible than a static table, the deflationary view holds that they remain reliant on superficial $n$-gram statistics or, at best, a broader array of complex heuristics based on shallow correlations. They lack the sensitivity to underlying structure associated with genuine understanding.
@@
Yet recent findings in mechanistic interpretability (MI)—the emerging field endeavoring to reverse-engineer the internal computations of these models—render this picture increasingly untenable. LLMs may be trained to perform next-token prediction, but the training objective tells us little about how they fulfill that task (Goldstein and Levinstein, 2024; Levinstein and Herrmann, 2024; Herrmann and Levinstein, 2025); and, as the findings we synthesize here demonstrate, surprisingly sophisticated mechanisms can emerge in response to this deceptively simple objective.
@@
Some of these mechanisms might indeed lie at the lower end of the hierarchy, generating impressive behavior through relatively unsophisticated means. There is research suggesting that LLMs can memorize training fragments verbatim (Carlini et al., 2023; Aerni et al., 2024; Stoehr et al., 2024), detect $n$-grams and learn $n$-gram statistics (Svete et al., 2024; Voita et al., 2024; Chang, 2025), and discover heuristics exploiting shallow correlations (see Du et al., 2023).
@@
A first indication that next-token prediction can drive the emergence of higher-level mechanisms is the phenomenon of “grokking,” which one researcher discovered after forgetting to stop a training run before leaving for vacation (Power, 2022; Power et al., 2022; Liu et al., 2022b, a; Nanda et al., 2023a; Varma et al., 2023). Grokking—a reference to the Heinleinian term for an alien form of “deep understanding”—describes a sudden shift during training where a model, after a long period of seemingly doing nothing but memorizing its training data, abruptly transitions to strong generalization on unseen data. Interestingly, this transition is typically accompanied by a decrease in the model’s internal complexity (measured by its “Minimum Description Length” or other proxies): the model appears to discard its sprawling collection of memorized examples in favor of something more compressed and generalizable (DeMoss et al., 2024). This is exactly the kind of shift that the deflationary picture struggles to explain.
LLMs offer unparalleled access to the internal computations underlying such shifts. Unlike with biological brains, the training histories and internal operations of LLMs can be precisely and comprehensively measured, replicated, and manipulated. Yet the field of MI is still in what Kuhnians would describe as a “pre-paradigmatic” stage: there is a growing collection of intriguing empirical results, but little consensus on an overarching framework and key conceptual and methodological questions (see Sharkey et al., 2025; Olah et al., 2020). This makes the field particularly apt to benefit from the theoretical and conceptual work that philosophy can contribute (Williams et al., 2025; Ayonrinde and Jaburi, 2025). Indeed, it is only once the empirical findings from MI are integrated within a theoretical account of understanding that their full import for debates over AI understanding becomes apparent.
@@
We start from the assumption that understanding fundamentally involves apprehending relational structure—appreciating the connections binding together the various aspects of an object, situation, or subject matter. There is significant convergence on this broadly unificationist idea in the epistemological literature. Wittgenstein (1953, §122) described understanding as “seeing connections,” while Kvanvig (2018, p. 697) characterizes it as grasping the explanatory and conceptual links between items of information. Similar accounts describing understanding as the ability to see how parts relate or “hang together” have been advanced by Grimm (2011) and Riggs (2003); indeed, the term “comprehension” itself, rooted in the Latin com- (“together”) and prehendere (“to seize”), evokes the act of grasping disjointed parts into a unified whole.
On this view, understanding constitutes a cognitive achievement distinct from the mere accumulation of facts or the exercise of abstract, domain-general reasoning. It allows an agent to navigate novel circumstances and manage counterfactual predictions (Hills, 2015; De Regt, 2017). While this competence can be grounded in a grasp of logical, probabilistic, or conceptual relations (Kvanvig, 2003, 2018), the most vital connections are frequently explanatory (Baumberger et al., 2017; Khalifa, 2017), involving the identification of causal processes (Grimm, 2006; Pritchard, 2014) or unifying principles and laws (Friedman, 1974; Kitcher, 1981; Belkoniene, 2023).
@@ 2  Conceptual Understanding
How might a system trained solely on next-token prediction develop the capacity to form internal representations that enable it to recognize and track entities and properties across diverse textual descriptions? This is the central question addressed by our first proposed tier of understanding: conceptual understanding. Just as humans form concepts to make sense of the world and track entities or properties across different encounters, we propose that LLMs develop internal representations—“features,” in MI terminology—that serve a similar function: they enable an LLM to grasp the connections between diverse descriptions and manifestations of an entity or property, effectively “same-tracking” it (Millikan, 2017) by subsuming its different presentations under a single, unifying representation.
@@ 2.1 The Emergence of Features
Using techniques such as sparse autoencoders (which we explain below), MI research is increasingly uncovering direct evidence of such features within the internal workings of LLMs. While early studies often focused on smaller toy models (Bricken et al., 2023), recent work has successfully identified features in state-of-the-art models. Indeed, Templeton et al. (2024) detail precisely such a “Golden Gate Bridge feature” in Claude 3 Sonnet. It activates reliably in response to various textual references to the bridge, such as “the colored bridge in SF,” and it does so across different languages and even in response to images.
@@
To sustain comparison with human concepts—at least on some theories of concepts—the Golden Gate Bridge feature needs to meet additional criteria. Drawing on Functional Role Semantics (Field, 1977; Harman, 1987; Block, 1987; Peacocke, 1992), recent work has formulated criteria for genuine representation in LLMs. Yetman (2025) and Williams (2026) argue that to count as a representation of X, a feature must play a causal role in the mechanisms generating the LLM’s robust X-related behavior and must be exploited in a way that explains successful task performance.
@@
Templeton et al. (2024) tested for the causal efficacy of the Golden Gate Bridge feature by intervening on it using a technique called “steering” (see also Bayat et al., 2025). [^4] Readers can experiment with steering for themselves at: https://www.neuronpedia.org/. They observed that amplifying the activation of the “Golden Gate Bridge” feature causes Claude 3 Sonnet to fixate on the bridge and consistently guide conversations on any topic back to it. When the activation of the Golden Gate Bridge feature is “clamped to ten times its max” (i.e. forced to take on a fixed value of ten times its maximum under normal operation), Claude ceased to answer the question “What is your physical form?” with: “I don’t actually have a physical form. I am an artificial intelligence,” and instead declared: “I am the Golden Gate Bridge. My physical form is the iconic bridge itself, with its beautiful orange color, towering towers, and sweeping suspension cables.” [^5] Anthropic temporarily allowed users to chat with such a steered model and called it “Golden Gate Claude” (Anthropic, 2024). The example above emerged from that experiment.
@@
Such sensitivity to intervention suggests that the feature is more than a passive correlate: it is a causal lever the system actively exploits to generate its output. This suggests the model has acquired a functional analogue to what we call “conceptual understanding”: it has learned to map diverse inputs to a single, unifying, and causally potent internal representation. And once we know that a feature plays this causal role, we are also in a position to turn off the feature, which is significant for AI safety and alignment (e.g., Karvonen and Marks, 2025).
@@
The fact that steering results in Claude misapplying the term “Golden Gate Bridge” to describe its own physical form also evokes another criterion for concept possession: that bona fide concepts are normative in allowing for a distinction between correct and incorrect applications. [^6] For a survey of the literature on the normativity of concepts, see Ginsborg (2018). This distinction needs to be grounded in something sufficiently independent from what seems correct to the individual concept-user. Some hold that this merely requires being subject to normative assessment by third parties (Hlobil, 2015)—a condition LLMs plausibly meet, both in their training and once deployed. Others maintain that normativity enters the picture once a certain internal state is imbued with the function of representing X (Wright, 1973; Millikan, 1989; Neander, 2017). If a state has the function of representing X, then activating it for Y counts as an error. As Mollo and Millière (2025) point out, LLMs meet the normativity condition on this account as well, because functions are acquired through a history of selection for certain effects, and the training process of LLMs is a form of selection through learning. They argue that this stabilizes the functions of the models’ internal states in the right way to warrant talk of representations that can be correctly or incorrectly applied.
@@
There are, then, several significant respects in which the Golden Gate Bridge feature sustains comparison with human concepts. And Templeton et al. (2024) identified millions of other features in Claude 3 Sonnet. Many corresponded to fairly abstract concepts, such as transit infrastructure—a feature which activated in response to descriptions or images of trains, ferries, tunnels, bridges, and even wormholes. Other examples included features corresponding to famous individuals and specific countries, to properties of code, to emotional expressions, and to slurs. Features seem to get activated very robustly, even in cases where one might not expect it: the eye feature identified by Tarng et al. (2025) activates not only on textual descriptions of eyes across multiple languages, but also on the specific characters used to represent eyes in ASCII-art faces (such as the colons in ":)" or the carets in "^_^") and even on lines of SVG code that render eyes when executed as images. [^7] Examples from other studies relying on linear probing include space and time features in Llama-2—such as longitude and latitude features and features for specific historical periods (Gurnee and Tegmark, 2024)—as well as sentiment features (Tigges et al., 2024).
@@
Such features are not explicitly programmed, but emerge as a consequence of training the model to meet its objective: next-token prediction. To excel at this task, an LLM must learn to differentiate inputs effectively (see Beckmann, 2025, §4.2). The “Golden Gate Bridge” feature allows the model to correctly generate upcoming tokens when responding to the question “What’s the orange bridge in SF called?,” for example. More generally, if distinguishing between contexts discussing the Golden Gate Bridge and contexts discussing the Brooklyn Bridge helps predict the next token more accurately, the model will learn to draw that distinction. It does so by becoming sensitive to patterns in the input data that are statistically correlated with certain continuations. A feature, in this MI context, thus corresponds to a task-relevant dimension along which the model learns to differentiate inputs. The model learns that by attending to certain features of the input, it can reduce the penalty it receives for prediction errors (i.e. minimize its “loss function”).
@@
Because these differentiating dimensions are specifically learned with a view to predicting the correct output, however, the same features the model uses to differentiate inputs must also help it differentiate between more or less suitable outputs. How the model organizes its internal processing of inputs (input-differentiation in latent space) is thus ultimately a matter of what best supports the process of output-differentiation, because the whole learning process is guided by the goal of producing the correct output. If learning to distinguish the Golden Gate Bridge from the Brooklyn Bridge serves the model well in predicting the tokens that follow sequences such as “a famous bridge in America,” the model will learn the distinction. But if the distinction never pays off at the level of next-token prediction, the model will never learn it. Input-differentiation is wholly subservient to output-differentiation. In this respect, LLMs are the perfect embodiment of the American pragmatists’ view of concepts as subject to “differential testing by consequences” (Dewey, 1929, 286). “The test of ideas,” John Dewey wrote, “is found in the consequences of the acts to which the ideas lead” (1929, 136). For LLMs, the test of features is found in the consequences of the predictions to which the features lead.
@@
It is also important to note that not all features correspond to external entities or their properties. LLMs also develop features related to the task of text generation itself. Examples include a “developing a weapon” feature (Templeton et al., 2024), a more general “refusal to answer” feature (Arditi et al., 2024), as well as features for understanding summarization requests or detecting jokes (Minder et al., 2025). Some research even points to meta-cognitive features, such as features for the model’s confidence in its knowledge about an entity or in its answer to a question (Ferrando et al., 2025; Lindsey et al., 2025). The pressure to accurately predict upcoming tokens compels the model to learn an extremely wide array of features, corresponding to all the discovered patterns that prove instrumental to performing its task of next-token prediction.

### v5

@@ 1  Introduction
The striking competence of Large Language Models (LLMs) raises an intriguing question: are they just mimicking human intelligence by relying on superficial statistics, or do they form internal structures that are sufficiently sophisticated and specific to sustain comparisons with human understanding? [^1] The issue of understanding in LLMs is distinct from that of understanding with LLMs (Sullivan,, 2022; Beisbart,, 2025), though the two are linked: as Sullivan effectively argues, whether we can understand with LLMs depends notably on how much understanding there is in LLMs.
@@
The most parsimonious explanation is that LLMs possess no understanding whatsoever. This soberly deflationary view receives succor from the surprisingly rudimentary principle upon which LLMs are built. At its foundation, an LLM is trained on a simple task: next-token prediction. The model is given a sequence of text and tasked with predicting the most probable next word or sub-word unit (“token”). According to the deflationary view, LLMs become good at this task by relying entirely on superficial statistical patterns within the distribution of words in the training data. This suggests that LLMs still operate fundamentally like traditional $n$-gram models, only at an unprecedented scale: by tallying how frequently different sequences of $n$ words co-occur (e.g. how often “The best pet is a” is followed by “dog” versus “cat,” etc.), they compute the probability of a word following a given sequence of “n1n-1” words. On this view, LLMs are doing nothing more than learning and then matching statistical patterns derived from extensive word-counting; they lack any grasp of the concepts, facts, and principles underlying these patterns. [^2] See, e.g., Bender and Koller, (2020); Bender et al., (2021); Mondorf and Plank, (2024); Merrill et al., (2021).
@@
This picture effectively casts LLMs as real-world approximations of a “Blockhead”—Ned Block’s (1981) hypothetical system that mimics intelligence by memorizing a massive table of input–output pairs. [^3] Millière and Buckner, (2024) discuss the relationship between LLMs and Blockhead in detail. While LLMs may be more flexible than a static table, the deflationary view holds that they remain reliant on superficial $n$-gram statistics or, at best, a broader array of complex heuristics based on shallow correlations. They lack the sensitivity to underlying structure associated with genuine understanding.
@@
Yet recent findings in mechanistic interpretability (MI)—the emerging field endeavoring to reverse-engineer the internal computations of these models—render this picture increasingly untenable. LLMs may be trained to perform next-token prediction, but the training objective tells us little about how they fulfill that task (Goldstein and Levinstein,, 2024; Levinstein and Herrmann,, 2024; Herrmann and Levinstein,, 2025); and, as the findings we synthesize here demonstrate, surprisingly sophisticated mechanisms can emerge in response to this deceptively simple objective.
@@
Some of these mechanisms might indeed lie at the lower end of the hierarchy, generating impressive behavior through relatively unsophisticated means. There is research suggesting that LLMs can memorize training fragments verbatim (Carlini et al.,, 2023; Aerni et al.,, 2024; Stoehr et al.,, 2024), detect $n$-grams and learn $n$-gram statistics (Svete et al.,, 2024; Voita et al.,, 2024; Chang,, 2025), and discover heuristics exploiting shallow correlations (see Du et al.,, 2023).
@@
A first indication that next-token prediction can drive the emergence of higher-level mechanisms is the phenomenon of “grokking,” which one researcher discovered after forgetting to stop a training run before leaving for vacation (Power,, 2022; Power et al.,, 2022; Liu et al., 2022b, ; Liu et al., 2022a, ; Nanda et al., 2023a, ; Varma et al.,, 2023). Grokking—a reference to the Heinleinian term for an alien form of “deep understanding”—describes a sudden shift during training where a model, after a long period of seemingly doing nothing but memorizing its training data, abruptly transitions to strong generalization on unseen data. Interestingly, this transition is typically accompanied by a decrease in the model’s internal complexity (measured by its “Minimum Description Length” or other proxies): the model appears to discard its sprawling collection of memorized examples in favor of something more compressed and generalizable (DeMoss et al.,, 2024). This is exactly the kind of shift that the deflationary picture struggles to explain.
LLMs offer unparalleled access to the internal computations underlying such shifts. Unlike with biological brains, the training histories and internal operations of LLMs can be precisely and comprehensively measured, replicated, and manipulated. Yet the field of MI is still in what Kuhnians would describe as a “pre-paradigmatic” stage: there is a growing collection of intriguing empirical results, but little consensus on an overarching framework and key conceptual and methodological questions (see Sharkey et al.,, 2025; Olah et al.,, 2020). This makes the field particularly apt to benefit from the theoretical and conceptual work that philosophy can contribute (Williams et al.,, 2025; Ayonrinde and Jaburi,, 2025). Indeed, it is only once the empirical findings from MI are integrated within a theoretical account of understanding that their full import for debates over AI understanding becomes apparent.
@@
We start from the assumption that understanding fundamentally involves apprehending relational structure—appreciating the connections binding together the various aspects of an object, situation, or subject matter. There is significant convergence on this broadly unificationist idea in the epistemological literature. Wittgenstein (1953, §122) described understanding as “seeing connections,” while Kvanvig (2018, p. 697) characterizes it as grasping the explanatory and conceptual links between items of information. Similar accounts describing understanding as the ability to see how parts relate or “hang together” have been advanced by Grimm, (2011) and Riggs, (2003); indeed, the term “comprehension” itself, rooted in the Latin com- (“together”) and prehendere (“to seize”), evokes the act of grasping disjointed parts into a unified whole.
On this view, understanding constitutes a cognitive achievement distinct from the mere accumulation of facts or the exercise of abstract, domain-general reasoning. It allows an agent to navigate novel circumstances and manage counterfactual predictions (Hills,, 2015; De Regt,, 2017). While this competence can be grounded in a grasp of logical, probabilistic, or conceptual relations (Kvanvig,, 2003, 2018), the most vital connections are frequently explanatory (Baumberger et al.,, 2017; Khalifa,, 2017), involving the identification of causal processes (Grimm,, 2006; Pritchard,, 2014) or unifying principles and laws (Friedman,, 1974; Kitcher,, 1981; Belkoniene,, 2023).
@@ 2  Conceptual Understanding
How might a system trained solely on next-token prediction develop the capacity to form internal representations that enable it to recognize and track entities and properties across diverse textual descriptions? This is the central question addressed by our first proposed tier of understanding: conceptual understanding. Just as humans form concepts to make sense of the world and track entities or properties across different encounters, we propose that LLMs develop internal representations—“features,” in MI terminology—that serve a similar function: they enable an LLM to grasp the connections between diverse descriptions and manifestations of an entity or property, effectively “same-tracking” it (Millikan,, 2017) by subsuming its different presentations under a single, unifying representation.
@@ 2.1 The Emergence of Features
Using techniques such as sparse autoencoders (which we explain below), MI research is increasingly uncovering direct evidence of such features within the internal workings of LLMs. While early studies often focused on smaller toy models (Bricken et al.,, 2023), recent work has successfully identified features in state-of-the-art models. Indeed, Templeton et al. (2024) detail precisely such a “Golden Gate Bridge feature” in Claude 3 Sonnet. It activates reliably in response to various textual references to the bridge, such as “the colored bridge in SF,” and it does so across different languages and even in response to images.
@@
To sustain comparison with human concepts—at least on some theories of concepts—the Golden Gate Bridge feature needs to meet additional criteria. Drawing on Functional Role Semantics (Field,, 1977; Harman,, 1987; Block,, 1987; Peacocke,, 1992), recent work has formulated criteria for genuine representation in LLMs. Yetman, (2025) and Williams, (2026) argue that to count as a representation of X, a feature must play a causal role in the mechanisms generating the LLM’s robust X-related behavior and must be exploited in a way that explains successful task performance.
@@
Templeton et al., (2024) tested for the causal efficacy of the Golden Gate Bridge feature by intervening on it using a technique called “steering” (see also Bayat et al.,, 2025). [^4] Readers can experiment with steering for themselves at: https://www.neuronpedia.org/. They observed that amplifying the activation of the “Golden Gate Bridge” feature causes Claude 3 Sonnet to fixate on the bridge and consistently guide conversations on any topic back to it. When the activation of the Golden Gate Bridge feature is “clamped to ten times its max” (i.e. forced to take on a fixed value of ten times its maximum under normal operation), Claude ceased to answer the question “What is your physical form?” with: “I don’t actually have a physical form. I am an artificial intelligence,” and instead declared: “I am the Golden Gate Bridge. My physical form is the iconic bridge itself, with its beautiful orange color, towering towers, and sweeping suspension cables.” [^5] Anthropic temporarily allowed users to chat with such a steered model and called it “Golden Gate Claude” (Anthropic,, 2024). The example above emerged from that experiment.
@@
Such sensitivity to intervention demonstrates that the feature is more than a passive correlate: it is a causal lever that the system actively exploits to generate its output. This suggests that the model has acquired a functional analogue to what we call “conceptual understanding”: it has learned to map diverse inputs to a single, unifying, and causally efficacious internal representation. And once we know that a feature plays this causal role, we are also in a position to turn off the feature, which is significant for AI safety and alignment (e.g., Karvonen and Marks, 2025).
@@
The fact that steering results in Claude misapplying the term “Golden Gate Bridge” to describe its own physical form also evokes another criterion for concept possession: that bona fide concepts are normative in allowing for a distinction between correct and incorrect applications. [^6] For a survey of the literature on the normativity of concepts, see Ginsborg, (2018). This distinction needs to be grounded in something sufficiently independent from what seems correct to the individual concept-user. Some hold that this merely requires being subject to normative assessment by third parties (Hlobil,, 2015)—a condition LLMs plausibly meet, both in their training and once deployed. Others maintain that normativity enters the picture once a certain internal state is imbued with the function of representing X (Wright,, 1973; Millikan,, 1989; Neander,, 2017). If a state has the function of representing X, then activating it for Y counts as an error. As Mollo and Millière, (2025) point out, LLMs meet the normativity condition on this account as well, because functions are acquired through a history of selection for certain effects, and the training process of LLMs is a form of selection through learning. They argue that this stabilizes the functions of the models’ internal states in the right way to warrant talk of representations that can be correctly or incorrectly applied.
@@
There are, then, several significant respects in which the Golden Gate Bridge feature sustains comparison with human concepts. And Templeton et al. (2024) identified millions of other features in Claude 3 Sonnet. Many corresponded to fairly abstract concepts, such as transit infrastructure—a feature which activated in response to descriptions or images of trains, ferries, tunnels, bridges, and even wormholes. Other examples included features corresponding to famous individuals and specific countries, to properties of code, to emotional expressions, and to slurs. Features seem to get activated very robustly, even in cases where one might not expect it: the eye feature identified by Tarng et al., (2025) activates not only on textual descriptions of eyes across multiple languages, but also on the specific characters used to represent eyes in ASCII-art faces (such as the colons in ":)" or the carets in "^_^") and even on lines of SVG code that render eyes when executed as images. [^7] Examples from other studies relying on linear probing include space and time features in Llama-2—such as longitude and latitude features and features for specific historical periods (Gurnee and Tegmark,, 2024)—as well as sentiment features (Tigges et al.,, 2024).
@@
Such features are not explicitly programmed, but emerge as a consequence of training the model to meet its objective: next-token prediction. To excel at this task, an LLM must learn to differentiate inputs effectively (see Beckmann,, 2025, §4.2). The “Golden Gate Bridge” feature allows the model to correctly generate upcoming tokens when responding to the question “What’s the orange bridge in SF called?,” for example. More generally, if distinguishing between contexts discussing the Golden Gate Bridge and contexts discussing the Brooklyn Bridge helps predict the next token more accurately, the model will learn to draw that distinction. It does so by becoming sensitive to patterns in the input data that are statistically correlated with certain continuations. A feature, in this MI context, thus corresponds to a task-relevant dimension along which the model learns to differentiate inputs. The model learns that by attending to certain features of the input, it can reduce the penalty it receives for prediction errors (i.e. minimize its “loss function”).
@@
Because these differentiating dimensions are specifically learned with a view to predicting the correct output, however, the same features the model uses to differentiate inputs must also help it differentiate between more or less suitable outputs. How the model organizes its internal processing of inputs (input-differentiation in latent space) is thus ultimately a matter of what best supports the process of output-differentiation, because the whole learning process is guided by the goal of producing the correct output. If learning to distinguish the Golden Gate Bridge from the Brooklyn Bridge serves the model well in predicting the tokens that follow sequences such as “a famous bridge in America,” the model will learn the distinction. But if the distinction never pays off at the level of next-token prediction, the model will never learn it. Input-differentiation is wholly subservient to output-differentiation. In this respect, LLMs are the perfect embodiment of the American pragmatists’ view of concepts as subject to “differential testing by consequences” (Dewey,, 1929, 286). “The test of ideas,” John Dewey wrote, “is found in the consequences of the acts to which the ideas lead” (1929, 136). For LLMs, the test of features is found in the consequences of the predictions to which the features lead.
@@
It is also important to note that not all features correspond to external entities or their properties. LLMs also develop features related to the task of text generation itself. Examples include a “developing a weapon” feature (Templeton et al.,, 2024), a more general “refusal to answer” feature (Arditi et al.,, 2024), as well as features for understanding summarization requests or detecting jokes (Minder et al.,, 2025). Some research even points to meta-cognitive features, such as features for the model’s confidence in its knowledge about an entity or in its answer to a question (Ferrando et al.,, 2025; Lindsey et al.,, 2025). The pressure to accurately predict upcoming tokens compels the model to learn an extremely wide array of features, corresponding to all the discovered patterns that prove instrumental to performing its task of next-token prediction.